OpenStudy (maheshmeghwal9):

Given: - \[f(x)=\sin x - \cos x.\]Prove that : -\[f(1) > 0.\]

OpenStudy (anonymous):

\[f(1)=\sin 1 -\cos 1\]prove that \(\sin 1>\cos 1\) using the fact that \(1>\frac{\pi}{4}\)

OpenStudy (anonymous):

hint\[\sin1 -\cos 1=\cos1 . (\tan 1-1)\]

OpenStudy (anonymous):

write sin 1 = sin ( inverse sin (pi /2) ) and cos(1) = cos (inverse cos (0))

OpenStudy (anonymous):

your answer would be pi/2 - 0

OpenStudy (anonymous):

which is pi/2

OpenStudy (anonymous):

and is >0

OpenStudy (maheshmeghwal9):

Thank u:D